rebeccaxhawaii (rebeccaxhawaii):

hey i have an Algebra 2 question

OpenStudy (jango_in_dtown):

bring it on

rebeccaxhawaii (rebeccaxhawaii):

If a football player passes a football from 4 feet off the ground with an initial velocity of 36 feet per second, how long will it take the football to hit the ground? Use the equation h = -16t2 + 36t + 4

OpenStudy (jango_in_dtown):

h = -16t² + 36t + 4. When h= 0, we have 0= -16t² + 36t + 4. or 16t² − 36t − 4.=0 or 4t² − 9t − 1.=0 →t=[ − (−9)±√{(−9)²− 4×4×(− 1)}]/(2×4) =[ 9 ±√{81+ 16}]/(8) =[ 9±√{97}]/(8) =2. 356 Seconds

OpenStudy (jango_in_dtown):

it will hit the ground when h=0

OpenStudy (jango_in_dtown):

from it we caluculate for t, with the help of the given equation/ you will get two values: one positive and another one negative, but time cannot be negative. so we will take the positive value

rebeccaxhawaii (rebeccaxhawaii):

so h = -16t2 + 36t + 4= 0 is in standard form by converting it into a quadratic equation will give me my answer right ?

OpenStudy (jango_in_dtown):

yeah right by putting h=0

rebeccaxhawaii (rebeccaxhawaii):

k thanks one more thing i was trying to solve this but i didnt get the right answer

OpenStudy (jango_in_dtown):

ok tell me

rebeccaxhawaii (rebeccaxhawaii):

h = -16t2 + 36t + 4

rebeccaxhawaii (rebeccaxhawaii):

can you use the equation button by the typing box pllllssssssssssss when explaining it makes it easier to comprehend

OpenStudy (jango_in_dtown):

do you want me to solve it again? i mean the quadratic equation?

rebeccaxhawaii (rebeccaxhawaii):

yes pls

OpenStudy (jango_in_dtown):

we have \[-16t^2+36t+4=0\] divide throughout by -4 and you have \[4t^2-9t-1=0\]

OpenStudy (jango_in_dtown):

now you solve this quadratic

rebeccaxhawaii (rebeccaxhawaii):

no wait sorry

rebeccaxhawaii (rebeccaxhawaii):

3x2 = 7x - 3

rebeccaxhawaii (rebeccaxhawaii):

3^2=7x-3

OpenStudy (jango_in_dtown):

\[3x^2=7x-3\] implies \[3x^2-7x+3=0\]

rebeccaxhawaii (rebeccaxhawaii):

can u solve it though

OpenStudy (jango_in_dtown):

yeah obviously. The general quadratic form is \[ax^2+bx+c=0\] where a, b,c are constants. and the solution is given by \[x=(-b \pm \sqrt{b^2-4ac})/2a\]

rebeccaxhawaii (rebeccaxhawaii):

mhmm

OpenStudy (jango_in_dtown):

here a=3,b=-7,c=3 now plug in the values of a,b,c in the above formula

OpenStudy (jango_in_dtown):

the answer is \[x=(7\pm \sqrt{13})/6\]

OpenStudy (jango_in_dtown):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

howd you get 13 i got 43

OpenStudy (jango_in_dtown):

\[(-7)^2-4*3*3=49-36=13\]

rebeccaxhawaii (rebeccaxhawaii):

oh duh i forgot about the exponent thanks

OpenStudy (jango_in_dtown):

n.p.

OpenStudy (jango_in_dtown):

ok bye. @rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

WAIT

rebeccaxhawaii (rebeccaxhawaii):

@jango_IN_DTOWN

OpenStudy (jango_in_dtown):

ok.. but tell me fast

rebeccaxhawaii (rebeccaxhawaii):

what did you get when solving the football player one

OpenStudy (jango_in_dtown):

2.356secs approx

rebeccaxhawaii (rebeccaxhawaii):

no no just in quadratic form what was the form

OpenStudy (jango_in_dtown):

\[4t^2-9t-1=0\]

rebeccaxhawaii (rebeccaxhawaii):

no the quadradic formula

OpenStudy (jango_in_dtown):

Which part I cant get you

rebeccaxhawaii (rebeccaxhawaii):

what is the equation in quadratic form

OpenStudy (jango_in_dtown):

-16t² + 36t + 4.=0

rebeccaxhawaii (rebeccaxhawaii):

no thats standard

OpenStudy (jango_in_dtown):

h = -16t² + 36t + 4. When h= 0, we have 0= -16t² + 36t + 4. or 16t² − 36t − 4.=0 or 4t² − 9t − 1.=0 →t=[ − (−9)±√{(−9)²− 4×4×(− 1)}]/(2×4) =[ 9 ±√{81+ 16}]/(8) =[ 9±√{97}]/(8) =2. 356 Seconds

rebeccaxhawaii (rebeccaxhawaii):

no i just need the equation that is in standard form converted into quadratic form

OpenStudy (jango_in_dtown):

ohkar ax^2+bx+c=0

rebeccaxhawaii (rebeccaxhawaii):

that is standard i need it in quadratic

OpenStudy (jango_in_dtown):

\[x=(-b \pm \sqrt{b^2-4ac})/2a\]

rebeccaxhawaii (rebeccaxhawaii):

yes what is the equation when converted to that form

OpenStudy (jango_in_dtown):

it is the solution of the standard form

rebeccaxhawaii (rebeccaxhawaii):

no can you convert -16t² + 36t + 4=0 into quadratic form

OpenStudy (jango_in_dtown):

ok divide throught y -4 and you have \[4t^2-9t-1=0\]

OpenStudy (jango_in_dtown):

\[t={-(-9)\pm \sqrt{(-9)^2-4*4*(-1)}}/(2*4)\]

rebeccaxhawaii (rebeccaxhawaii):

why is it like that now ?

rebeccaxhawaii (rebeccaxhawaii):

it was postitves before

OpenStudy (jango_in_dtown):

here b=-9 which is negative

rebeccaxhawaii (rebeccaxhawaii):

but why did it change ?

OpenStudy (jango_in_dtown):

we divided by -4.. so a negative came

rebeccaxhawaii (rebeccaxhawaii):

thank you so much @jango_IN_DTOWN